A truck and a trailer weigh 4000 kg EACH. In order to accelerate at 2 m/s2, how much force would need to be produced?


1
Expert's answer
2021-01-26T08:41:31-0500

"F=ma\to F=(4000+4000)\cdot2=16000(N)" . Answer